# Data Analysis Prediction over Multiple Unseen Datasets: A Vector Embedding Approach

### v1

@@ Abstract
The massive increase in the data volume and dataset availability for analysts compels researchers to focus on data content and select high-quality datasets to enhance the performance of analytics operators. While selecting the highest quality data for analysis highly increases task accuracy and efficiency, it is still a hard task, especially when the number of available inputs is very large. To address this issue, we propose a novel methodology that infers the outcome of analytics operators by creating a model from datasets similar to the queried one. Dataset similarity is performed via projecting each dataset to a vector embedding representation. The vectorization process is performed using our proposed deep learning model NumTabData2Vec, which takes a whole dataset and projects it into a lower vector embedding representation space. Through experimental evaluation, we compare the prediction performance and the execution time of our framework to another state-of-the-art modelling operator framework, illustrating that our approach predicts analytics outcomes accurately. Furthermore, our vectorization model can project different real-world scenarios to a lower vector embedding representation and distinguish between them.
@@ III-C Dataset Selection
1:Data Lake Vectors $\mathbf{V}=\{\mathbf{v}_{1},\mathbf{v}_{2},\dots,\mathbf{v}_{n}\}$, $D_{o}$ vector $v_{o}$, range of clusters $S=(2,\dots,p)$
2:Initialize the number of cluster $s$ using Silhouette score: $s=SilhouetteScore(V,S)$
3:Initialize the $s$ cluster centroids $\mathbf{C}=\{\mathbf{c}_{1},\mathbf{c}_{2},\dots,\mathbf{c}_{s}\}$ randomly from the vectors $\mathbf{V}$.
@@
6:     for each vector $\mathbf{v}_{i}\in\mathbf{V}$ do
7:         Assign $\mathbf{v}_{i}$ to the nearest centroid based on Euclidean distance:
Assign $\text{Assign }\mathbf{v}_{i}\text{ to cluster }j=\arg\min_{j}\|\mathbf{v}_{i}-\mathbf{c}_{j}\|^{2}$
@@
$\mathbf{c}_{j}=\frac{1}{|\{\mathbf{v}_{i}\in C_{j}\}|}\sum_{\mathbf{v}_{i}\in C_{j}}\mathbf{v}_{i}$
@@
15:Find in which cluster the vector $v_{o}$ belongs,     $c=\textbf{K-Means}(v_{o})$
@@
17:Return Datasets $D_{r}$
@@
Algorithm 1, shown how k-means clustering is created and how the most relevant datasets $D_{r}$ are selected based on the target dataset $D_{o}$. Using the most optimal $s$ value with the highest Silhouette score, K-means clustering is performed on the vector representations $V$ of each dataset, grouping datasets with same features into the same cluster. Next, the algorithm uses the vector $v_{o}$ of dataset $D_{o}$ to find the closest cluster centroid. All datasets in that cluster are defined as the relevant datasets $D_{r}$. These datasets are then used to model the analytics operator at hand.
@@ IV-C Evaluation Results
In the weather dataset, the SVM operator results from sub-figures 5(a) and 5(b) show that using $300$ vectors in the representation space consistently led to more accurate operator models across all similarity methods. Specifically, cosine similarity in combination with the $300$-dimensional vector embedding reduced the error rate in operator predictions, demonstrating that projecting datasets into this representation space and applying cosine similarity improves the prediction accuracy on the modelled operator. For the MLP classifier from sub-figures 5(c) and 5(d), the results illustrate that using vector embeddings of size $200$ and K-Means clustering produced the most accurate MLP classifier operators.
@@
$200$V Cosine MLP $\mathbf{14.29}$ $\mathbf{4.03}$ $1.03$ $\mathbf{22.8}$
$200$V SR-$0.4$ MLP $15.95$ $13.31$ $1.02$ $1.77$

### v2

@@ Abstract
The rapid growth in data volume and the increasing availability of datasets present analysts with the challenge of selecting high-quality data to optimize the performance of analytics operations. While choosing the best-quality data significantly improves accuracy and efficiency, it remains a complex task, particularly when dealing with a vast number of inputs. To tackle this challenge, we introduce a novel approach that predicts analytics outcomes by building a model based on datasets similar to the target dataset. Our method determines dataset similarity by mapping each dataset into a vector embedding representation. This transformation is achieved through our deep learning model, NumTabData2Vec, which processes entire datasets and projects them into a lower-dimensional vector space. Through experimental evaluation, we compare our framework’s prediction accuracy and execution time against a leading state-of-the-art modelling operator, demonstrating that our approach effectively predicts analytics outcomes. Additionally, our vectorization model can capture and differentiate diverse real-world scenarios by representing them in a compact vector space.
@@ III-C Dataset Selection
1:Vectors $\mathbf{Z}=\{\mathbf{z}_{1},\mathbf{z}_{2},\dots,\mathbf{z}_{n}\}$, $D_{o}$ vector $z_{o}$, range of clusters $S=(2,\dots,p)$, Maximum size of cluster $max_{s}$, Minimum size of cluster $min_{s}$
2:Initialize the number of cluster $s$ using Silhouette score: $s=SilhouetteScore(Z,S)$
3:Initialize the $s$ cluster centroids $\mathbf{C}=\{\mathbf{c}_{1},\mathbf{c}_{2},\dots,\mathbf{c}_{s}\}$ randomly from the vectors $\mathbf{Z}$.
@@
6:     for each vector $\mathbf{z}_{i}\in\mathbf{Z}$ do
7:         Assign $\mathbf{z}_{i}$ to the nearest centroid based on Euclidean distance:
Assign $\text{Assign }\mathbf{z}_{i}\text{ to cluster }j=\arg\min_{j}\|\mathbf{z}_{i}-\mathbf{z}_{j}\|^{2}$
@@
$\mathbf{c}_{j}=\frac{1}{|\{\mathbf{z}_{i}\in C_{j}\}|}\sum_{\mathbf{z}_{i}\in C_{j}}\mathbf{v}_{i}$
@@
15:Find in which cluster the vector $z$ belongs, $c=\textbf{K-Means}(z)$
@@
17:Check the number of datasets in $D_{r}$ and update it if it does not meet the $min_{s}$ and $max_{s}$.
18:Return Datasets $D_{r}$
@@
Algorithm 1, shown how k-means clustering is created and how the most relevant datasets $D_{r}$ are selected based on the target dataset $D_{o}$. Using the most optimal $s$ value with the highest Silhouette score, K-means clustering is performed on the vector representations $V$ of each dataset, grouping datasets with same features into the same cluster. Next, the algorithm uses the vector $v_{o}$ of dataset $D_{o}$ to find the closest cluster centroid. All datasets in that cluster are defined as the relevant datasets $D_{r}$. These datasets are then used to model the analytics operator at hand. However, we defined a maximum and minimum size for $D_{r}$, and if these conditions are not met, datasets are either removed from the cluster or added based on their distance from the cluster centroid. These small adjustments are only made in cases where the clustering technique does not yield results that satisfy our requirements.
@@ IV-C Evaluation Results
In the weather dataset, the SVM operator results from sub-figures 5(a) and 5(b) show that using $300$ vectors in the representation space consistently led to more accurate operator models across all similarity methods. Specifically, cosine similarity in combination with the $300$-dimensional vector embedding reduced the error rate in operator predictions, demonstrating that projecting datasets into this representation space and applying cosine similarity improves the prediction accuracy on the modelled operator. For the MLP classifier (Figures 5(c) and 5(d)), the results illustrate that using vector embeddings of size $300$ and Cosine similarity- produced the most accurate MLP classifier operators.
@@
$300$V Cosine MLP $\mathbf{14.29}$ $\mathbf{4.03}$ $1.03$ $\mathbf{22.8}$
$300$V SR-$0.4$ MLP $15.95$ $13.31$ $1.02$ $1.77$